# A Survey of Real-Time Support, Analysis, and Advancements in ROS 2

### v1

@@ 2.1 ROS 2
Figure 1: ROS 2, with its layers and interactions.
ROS 2 applications are realized through a distributed network of components. Each component is realized as a node, which is the fundamental unit of composition in ROS 2 systems and encapsulates a set of related functionalities. The functionalities of a node are implemented as callbacks, which are functions that are executed in response to specific events. These events are triggered by different task types, such as subscribers that receive messages (event-driven) or timers that elapse after a specified time interval (time-driven).
@@
Another mechanism to activate callbacks is via timers, which trigger the execution of a callback periodically, according to a specified time interval. Furthermore, ROS 2 also provides a mechanism for remote procedure calls (RPCs) through services. These feature a client-server architecture, where a client callback sends a request to a service server, which processes the request and returns a response, which is then processed by a client callback.
@@
Historical Context for Polling. As highlighted by Teper et al. [teper2025reconciling], the polling mechanism of ROS 2 is significantly different from classical scheduling theory of periodic tasks, which assumes every job will be executed. Indeed, for ROS 2 versions after “Dashing”, timer callbacks may be skipped if the executor detects that the time between a previous timestamp and the start time of the next timer job is longer than the period of the timer callback itself.
@@ 2.2 Beyond ROS 2
The choice of middleware has significant implications for system performance and real-time behavior. Because each implementation has its own internal architecture, threading model, and configuration parameters, a generic timing analysis is insufficient. For example, Sciangula et al. [sciangula2023bounding] provided a compositional model that generalizes beyond the specific implementation, taking into account the common aspects. Additionally, they provide a concrete instantiation of the model for FastDDS to devise a response-time analysis. Understanding and modeling the specific timing properties of the chosen middleware is crucial for accurately predicting end-to-end latencies and ensuring that real-time requirements are met.
@@
Figure 4 illustrates the timing behavior of DDS-based synchronous communication in ROS 2. Callback $C_{1}$ sends messages to callback $C_{2}$ through a buffer. In this scenario, three messages are sent from $C_{1}$ to $C_{2}$. Due to the polling mechanism of ROS 2, the messages are stored in the buffer and delivered to $C_{2}$ only at the next polling point. As a result, for the second and third message, while being available before the start of $C_{2,1}$ and $C_{2,2}$, respectively, they are only processed one processing window later due to the buffering mechanism.
@@
Moreover, this flow-controller thread can itself be configured with different scheduling policies, such as HIGH_PRIORITY (i.e., fixed-priority), FIFO, or ROUND-ROBIN. The choice between synchronous and asynchronous mode, combined with the scheduling policy of the flow-controller thread, can substantially impact the overall latency and determinism of message delivery in ROS 2 systems. Furthermore, on the receiver side, messages are typically delivered by a DDS listener thread to the ROS 2 executed in FIFO order, adding another stage to the communication pipeline that must be considered in a full system analysis.
@@ 2.3 Overview of the ROS 1 Approaches for Real Time
While the Linux kernel has since gained powerful real-time capabilities, e.g., through tools for its fine-grained configuration [de2022operating, de2025timerlat], the PREEMPT_RT patch [de2020thread, de2017automata], and the SCHED_DEADLINE scheduler [phoronix2024, Lelli2016], ROS 1 itself lacks the internal architecture and APIs needed to effectively leverage them. In contrast, the design of ROS 2 aims at explicitly targets real-time support through a modular execution model, DDS-based communication, and configurable DDS QoS policies [perez2015modeling]. The fundamental design limitation of ROS 1 prompted numerous research efforts to retrofit real-time capabilities onto the ROS 1 framework, which can be broadly categorized into two main strategies.
@@ 3 Analysis of Real-Time Behavior in ROS 2
We start the survey by summarizing the papers on the real-time analysis of ROS 2. Achieving predictable real-time performance in ROS 2 requires a deep understanding of the various factors that contribute to the real-time behavior. Over the past several years, a significant body of research has emerged to formally analyze these factors from different perspectives. This section provides a comprehensive overview of these analytical efforts. We begin by examining the core of ROS 2’s internal scheduling logic, covering works that model the behavior of both the single-threaded (Section˜3.1) and multi-threaded executors (Section˜3.2). We then explore studies that focus on the communication stack, discussing techniques for bounding delays introduced by lower-level middleware like DDS and specialized message synchronization filters in Section˜$3.$ Subsequently, we explore an alternative approach based on formal verification and model checking in Section˜3.4, which explores the verification of timing properties in a different way. Finally, a taxonomy is presented in Section˜3.5 to systematically classify and compare the different analytical approaches discussed throughout this section.
@@ 3.1 Analysis of the Single-Threaded Executor
The core of their analysis is a search for the worst-case scenario within a so-called “busy window”, which can be informally defined as a period of continuous processing. The interested reader in the correct and formal definition is left to the original paper. The response-time analysis presented in the paper builds on computing the response-time of arbitrary callback instances released at an arbitrary point in time AA within the busy window. The per-callback response time is then computed as the maximum among all possible instances released at arbitrary values of AA. Clearly, this requires bounding and discretizing the space of possible time instants AA.
@@
Analysis of Reaction Time and Data Age. In addition to the traditional response-time metrics, recent research has introduced a shift to metrics that focus on data propagation for ROS 2. This line of research was initiated by Teper et al. [Teper2022], who expanded the system model to cover more diverse communication patterns. Beyond the typical publish-publish mechanism (termed inter-node communication), their work also formalized intra-node communication, where callbacks exchange data by storing it within the same node, and external communication, where callbacks interact with external data interfaces. Critically, they focus on two new performance metrics:
@@ 3.3 Bounding Communication and Synchronization Delays
Lower-Level Middleware (DDS). In ROS 2, communication occurs in two main ways: intra-process and inter-process communication. Intra-process communication, which handles message exchanges within the same executor, is optimized to avoid unnecessary copies when sending messages from publishers to subscribers. In contrast, inter-process communication manages the more complex task of connecting publishers and subscribers across different executors, often on separate machines, and relying on the underlying DDS middleware. Because the performance of distributed robotic systems hinges on the efficiency of this layer, most real-time research has focused on analyzing the overheads of inter-process communication.
@@
The ApproximateTime policy is the most widely analyzed. It groups messages that arrive within a specified time window, tolerating a small time disparity. Research on this policy has evolved over several years. Li et al. [li2022worst] first modeled and analyzed this time disparity, initially for systems with unlimited buffers and later extending the analysis to more realistic limited-size buffers [li2023modeling]. Moreover, the ApproximateTime policy potentially needs messages to remain in the message synchronizer in order to work, introducing additional latencies. Li et al. [li2023modeling] were the first to formally model and experimentally measure these latencies, introducing the concepts of passing latency and reaction latency. The passing latency is defined as the time a message spends in the synchronizer before being output, while the reaction latency is the time from message arrival to output, including any necessary discarding of older messages. In 2023, Li et al. [li2023worst] provided a formal model and experimental measurements of these latencies. Later, Wu et al. [wu2024improving] built upon [li2023worst] to provide more accurate estimates. Separately, Sun et al. [sun2023seam] identified a limitation of the ApproximateTime policy: it requires prior knowledge of incoming message timing characteristics, and any inaccuracies in this knowledge can significantly degrade performance. To address this, they proposed a new message synchronizer called SEAM that does not rely on a priori information about the messages.
The LatestTime policy is a more recent addition to ROS 2, designed to maximize output frequency. It achieves this by implementing a zero-order-hold mechanism, storing the latest message from each topic and publishing the complete set whenever any single topic receives a new message. Wu et al. [wu2024modeling] were the first to model the latency of this policy and uncovered a critical flaw in its design that could lead to unbounded latencies under certain scenarios. Finally, the impact on real-time metrics of different sensor fusion patterns has been recently explored by Sobhani et al. [sobhani2025fusionpatterns] in the context of more abstractly modeled autonomous driving systems.
@@ 3.5 Taxonomy of Analysis Methods
  • Table˜2 (DDS Communication Analysis) focuses specifically on the papers that analyze the DDS middleware, as detailed in Section˜$3.$ This table highlights the communication mode, the specific DDS implementation targeted, and the primary focus of the analysis, from delay modeling to system-wide optimization.
  • Table˜3 (Message Synchronization Analysis) summarizes the research on message filters and synchronizers, discussed in Section˜$3.$ It categorizes papers by the synchronizer policy they analyze and key contributions, e.g., modeling time disparity or identifying design flaws.
@@
Table 2: Lower-level Middleware DDS Communication Analysis in ROS 2
@@
Sun et al. [sun2023seam] SEAM (Custom) Novel synchronizer New synchronizer without a priori knowledge for better robustness
@@ 4.1 Customized Executor Design
Very recently, Ishikawa et al. [ishikawa2025work] proposed the middleware-transparent ROS 2 executor, which establishes a persistent one-to-one mapping between callbacks and Linux threads, superseding the executor’s behavior and allowing callbacks to be executed directly by the OS’s scheduling policy. Finally, these principles of priority-driven scheduling have also been adapted to resource-constrained environments such as micro-ROS. The default executor in micro-ROS, known as the RCLC Executor, suffers from two critical issues: due to its FIFO-based data processing strategy, it is unaware of callback priority, leading to priority inversion during callback execution and data processing. Additionally, the executor enters a blocking state during data transmission, leading to inefficient CPU utilization and wasted resources. To solve these issues, Wang et al. [wang2024improving] proposed PoDS, a Priority-Driven chain-aware Scheduling system implemented with a new Timing-Deterministic and Efficient (TIDE) Executor. The TIDE Executor introduces two priority-aware queues: a timer queue sorted by timestamps and a ready queue sorted by priority. It also incorporates a priority-based data-processing mechanism and a parallel transmission model to minimize blocking and reduce priority inversion. These modifications make callback execution and data handling more deterministic and efficient. Experimental results confirm that PoDS significantly outperforms the default micro-ROS in terms of runtime predictability and real-time stability, making it a promising advancement for low-resource robotic systems.
@@ 4.2 System-Level Enhancements
Optimizing Inter-Process Communication. Beyond CPU scheduling, researchers have proposed system-level enhancements targeting the performance and predictability of ROS 2’s inter-process communication (IPC), often focusing on limitations within the default DDS middleware or exploring alternatives. Kim et al. [kim2025cros] identified the lack of priority propagation across ROS 2’s multi-layered architecture, leading to priority inversion where low-priority DDS or kernel threads delay high-priority application messages. To address this, they developed the Cros-Rt framework, which explicitly aligns thread priorities across the application, middleware, and kernel layers to ensure consistent communication across layers that uphold priority ordering. Orthogonally, Ishikawa et al. [ishikawaaso2025ros2agnocastsupporting] proposed the ROS 2 Agnocast framework to optimize the data transfer efficiency within ROS 2 by using zero-copy shared memory for IPC. They observed that current applications are often bottlenecked by data serialization and copying overheads in DDS, especially for large messages, but also identified key limitations in existing zero-copy DDS implementations, such as the inability to handle unsized message types (e.g., images or point clouds). They introduce a novel publish/subscribe solution through Agnocast using shared memory and pointer-swapping that supports unsized message types, enabling true zero-copy IPC and further enhancing communication efficiency through latency reduction and throughput improvement. Specific middleware protocols for large data objects have been investigated also by Peeck et al. [Peeck2021, Peeck2023] in the context of the DDS.
Enhancing GPU and Accelerator Management. ROS 2 serves as a powerful resource manager for CPU tasks, but it lacks built-in support for GPU resource management. Many critical functionalities in autonomous systems, such as perception and intelligent control, require intensive GPU computation. However, under default ROS 2, these components must submit their GPU workloads independently, without coordination.

### v2

@@ 2.1 ROS 2
ROS 2 applications are realized through a distributed network of components, as illustrated in LABEL:{fig:components}. Each component is realized as a node, which is the fundamental unit of composition in ROS 2 systems and encapsulates a set of related functionalities. The functionalities of a node are implemented as callbacks, which are functions that are executed in response to specific events. These events are triggered by different task types, such as subscribers that receive messages (event-driven) or timers that elapse after a specified time interval (time-driven).
Figure 2: ROS 2 components.
@@
Another mechanism to activate callbacks is via timers, which trigger the execution of a callback periodically, according to a specified time interval. Furthermore, ROS 2 also provides a mechanism for remote procedure calls (RPCs) through services. These feature a client-server architecture, where a client callback sends a request to a service server, which processes the request and returns a response, which is then processed by a client callback, as shown in LABEL:{fig:components}.
@@
Historical Context for Polling. As highlighted by Teper et al. [teper2025reconciling], the polling mechanism of ROS 2 is significantly different from classical scheduling theory of periodic tasks, which assumes every job will be executed. Indeed, for ROS 2 versions after “Dashing”, timer callbacks are be skipped if the executor detects that the time between the activation time of the timer callback’s previous job and the start time of its next job is longer than the period of the timer callback itself.
@@ 2.2 Beyond ROS 2
The choice of middleware has significant implications for system performance and real-time behavior. Because each implementation has its own internal architecture, threading model, and configuration parameters, a generic timing analysis is insufficient. For example, Sciangula et al. [sciangula2023bounding] provided a compositional model that generalizes beyond the specific implementation, taking into account the common aspects. Additionally, they provide a concrete instantiation of the model for FastDDS to devise a response-time analysis. This work is further extended in [Sciangula2025] to also support all the policies of FastDDS and to include multiple communication modes, as well as integrating the analysis with ROS 2. Understanding and modeling the specific timing properties of the chosen middleware is crucial for accurately predicting end-to-end latencies and ensuring that real-time requirements are met.
@@
We illustrate these behaviors using the example system in Figures 5 and 6, which consist of a callback $C_{1}$ that publishes three messages to a subscriber callback $C_{2}$ via DDS. Each callback is run by its own executor. We assume that the executor of $C_{2}$ starts after time point $6$, and that each job of callback $C_{1}$ publishes at the end of its execution.
Figure 5 illustrates the timing behavior of DDS-based synchronous communication in ROS 2. As soon as a job of callback $C_{1}$ finishes its publish operation, the message is guaranteed to be stored in the buffer of $C_{2}$. Since messages may arrive while the executor of $C_{2}$ is already processing a previous message, the buffer stores multiple messages that are then processed in FIFO order, one per processing window.
Figure 6 illustrates the timing behavior of DDS-based asynchronous communication for the same scenario. In contrast to the synchronous case, the publishing callback $C_{1}$ does not wait for the message to be sent. Instead, the message is handed off to the flow-controller thread, which handles the actual transmission. As a result, the execution time of each $C_{1}$ instance is shorter, since it no longer includes the time spent on the publish operation. However, the subscriber’s buffer is only updated once the flow-controller thread has completed the transmission, which may introduce additional latency compared to the synchronous mode. This is illustrated between time points $6$ and $8$, in which the buffer is empty, while message $m_{2}$ is still being processed by the flow-controller thread.
@@
Moreover, this flow-controller thread can itself be configured with different scheduling policies, such as HIGH_PRIORITY (i.e., fixed-priority), FIFO, or ROUND-ROBIN. The choice between synchronous and asynchronous mode, combined with the scheduling policy of the flow-controller thread, can substantially impact the overall latency and determinism of message delivery in ROS 2 systems. Furthermore, on the receiver side, messages are typically delivered by a DDS listener thread to the ROS 2 executor in FIFO order, adding another stage to the communication pipeline that must be considered in a full system analysis.
[…]
Figure 6: DDS-based asynchronous inter-node communication.
@@ 2.3 Overview of the ROS 1 Approaches for Real Time
While the Linux kernel has gained powerful real-time capabilities, e.g., through tools for its fine-grained configuration [de2022operating, de2025timerlat], the PREEMPT_RT patch [de2020thread, de2017automata], and the SCHED_DEADLINE scheduler [phoronix2024, Lelli2016], ROS 1 itself lacks the internal architecture and APIs needed to effectively leverage them. In contrast, the design of ROS 2 aims at explicitly targets real-time support through a modular execution model, DDS-based communication, and configurable DDS QoS policies [perez2015modeling]. The fundamental design limitation of ROS 1 prompted numerous research efforts to retrofit real-time capabilities onto the ROS 1 framework, which can be broadly categorized into two main strategies.
@@ 3 Analysis of Real-Time Behavior in ROS 2
We start the survey by summarizing the papers on the real-time analysis of ROS 2. Achieving predictable real-time performance in ROS 2 requires a deep understanding of the various factors that contribute to the real-time behavior. Over the past several years, a significant body of research has emerged to formally analyze these factors from different perspectives. This section provides a comprehensive overview of these analytical efforts. We begin by examining the core of ROS 2’s internal scheduling logic, covering works that model the behavior of both the single-threaded (Section˜3.1) and multi-threaded executors (Section˜3.2). We then explore studies that focus on the communication stack, discussing techniques for bounding delays introduced by lower-level middleware like DDS and specialized message synchronization filters in Section˜3.3. Subsequently, we explore an alternative approach based on formal verification and model checking in Section˜3.4, which explores the verification of timing properties in a different way. Finally, a taxonomy is presented in Section˜3.5 to systematically classify and compare the different analytical approaches discussed throughout this section.
@@ 3.1 Analysis of the Single-Threaded Executor
The core of their analysis is a search for the worst-case scenario within a so-called “busy window”, which can be informally defined as a period of continuous processing. The interested reader in the correct and formal definition is left to the original paper. The response-time analysis presented in the paper builds on computing the response-time of arbitrary callback instances released at an arbitrary point in time $A$ within the busy window. The per-callback response time is then computed as the maximum among all possible instances released at arbitrary values of $A$. Clearly, this requires bounding and discretizing the space of possible time instants $A$.
@@
Analysis of Reaction Time and Data Age. In addition to the traditional response-time metrics, recent research has introduced a shift to metrics that focus on data propagation for ROS 2. This line of research was initiated by Teper et al. [Teper2022], who expanded the system model to cover more diverse communication patterns. Beyond the typical publish-subscribe mechanism (termed inter-node communication), their work also formalized intra-node communication, where callbacks exchange data by storing it within the same node, and external communication, where callbacks interact with external data interfaces. Critically, they focus on two new performance metrics:
@@
In addition to the aforementioned methods, new research directions are emerging. For example, Lee et al. [lee2022probabilistically] and Han and Kim [han2023minimizing] propose probabilistic real-time analysis and optimization methods for end-to-end latency in autonomous-driving task graphs, using Autoware as a case study; however, these works target more general scheduling assumptions and are not specific to the ROS 2 scheduling policy.
@@ 3.3 Bounding Communication and Synchronization Delays
Intra-Process Communication and Lower-Level Middleware (DDS). In ROS 2, communication occurs in two main ways: intra-process and inter-process communication. Intra-process communication, which handles message exchanges within the same executor, is optimized to avoid unnecessary copies when sending messages from publishers to subscribers.
Intra-process communication has been recently characterized by Luo et al. [luo2025scalability], who showed that its overhead is not constant and can vary based on the message configuration, workload structure, and usage semantics. To improve scalability, the paper proposes guidelines for configuring intra-process communication and an allocation-free message pooling mechanism.
In contrast, inter-process communication manages the more complex task of connecting publishers and subscribers across different executors, often on separate machines, and relying on the underlying DDS middleware. Because the performance of distributed robotic systems hinges on the efficiency of this layer, most real-time research has focused on analyzing the overheads of inter-process communication.
@@
The ApproximateTime policy is the most widely analyzed. It groups messages that arrive within a specified time window, tolerating a small time disparity. Research on this policy has evolved over several years. Li et al. [li2022worst] first modeled and analyzed this time disparity, initially for systems with unlimited buffers and later extending the analysis to more realistic limited-size buffers [li2023modeling]. Moreover, the ApproximateTime policy potentially needs messages to remain in the message synchronizer in order to work, introducing additional latencies. Li et al. [li2023modeling] were the first to formally model and experimentally measure these latencies, introducing the concepts of passing latency and reaction latency. The passing latency is defined as the time a message spends in the synchronizer before being output, while the reaction latency is the time from message arrival to output, including any necessary discarding of older messages. In 2023, Li et al. [li2023worst] provided a formal model and experimental measurements of these latencies. Later, Wu et al. [wu2024improving] built upon [li2023worst] to provide more accurate estimates.
A notable limitation of the ApproximateTime policy is that it requires prior knowledge of incoming message timing characteristics, and any inaccuracies in this knowledge can significantly degrade performance [sun2023seam]. Two distinct approaches circumvent this limitation by not relying on a priori information about the messages, though they differ fundamentally in their design goals and guarantees.
The first is SEAM [sun2023seam], a custom message synchronizer that selects the earliest-arriving message from each input queue such that the time disparity among the selected messages remains within a user-specified threshold, thereby guaranteeing well-bounded time disparity without requiring any prediction of future message arrivals.
The second is the LatestTime policy, a more recent addition to ROS 2, which instead prioritizes maximizing output frequency. It achieves this by implementing a zero-order-hold mechanism, storing the latest message from each topic and publishing a complete set of the latest message per topic whenever any single topic receives a new message. However, unlike SEAM, LatestTime does not guarantee bounded time disparity. Wu et al. [wu2024modeling] were the first to model the latency of this policy and uncovered a critical flaw in its design that could lead to unbounded latencies under certain scenarios.
Finally, the impact on real-time metrics of different sensor fusion patterns has been recently explored by Sobhani et al. [sobhani2025fusionpatterns] in the context of more abstractly modeled autonomous driving systems. Here, the authors identified task types and fusion patterns, dependent on the number of inputs and the activation patterns, and evaluated their impact on the real-time performance of the system.
@@ 3.5 Taxonomy of Analysis Methods
  • Table˜2 (DDS Communication Analysis) focuses specifically on the papers that analyze the DDS middleware, as detailed in Section˜3.3. This table highlights the communication mode, the specific DDS implementation targeted, and the primary focus of the analysis, from delay modeling to system-wide optimization.
  • Table˜3 (Message Synchronization Analysis) summarizes the research on message filters and synchronizers, discussed in Section˜3.3. It categorizes papers by the synchronizer policy they analyze and key contributions, e.g., modeling time disparity or identifying design flaws.
@@
Table 2: Communication Analysis in ROS 2, either intra-process or DDS-based.
@@
Sciangula et al. [Sciangula2025] Sync+Async. inter-process FastDDS End-to-end latency analysis with FastDDS threads modeling, and integration with ROS 2 analysis
@@
Sun et al. [sun2023seam] SEAM (Custom) Novel synchronizer New synchronizer without a priori knowledge, guaranteeing well-bounded time disparity
@@ 4.1 Customized Executor Design
[…]
micro-ROS. Finally, these principles of priority-driven scheduling have also been adapted to resource-constrained environments such as micro-ROS. The default executor in micro-ROS, known as the RCLC Executor, suffers from two critical issues highlighted in [wang2023tide]: due to its FIFO-based data processing and batch-based callback scheduling, it may violate priority ordering, leading to priority inversion during callback execution and data processing. Additionally, the executor enters a blocking state during data transmission, leading to inefficient CPU utilization and wasted resources. To address these issues at the system level, Wang et al. [wang2024improving] proposed PoDS, a Priority-Driven chain-aware Scheduling system that integrates the Timing-Deterministic and Efficient (TIDE) executor introduced in [wang2023tide]. The TIDE Executor introduces two priority-aware queues: a timer queue sorted by timestamps and a ready queue sorted by priority. It also incorporates a priority-based data-processing mechanism and a parallel transmission model to minimize blocking and reduce priority inversion. These modifications make callback execution and data handling more deterministic and efficient. Experimental results in [wang2024improving] confirm that PoDS significantly outperforms the default micro-ROS in terms of runtime predictability and real-time stability, making it a promising advancement for low-resource robotic systems.
@@
Liu et al. [liu2025rosrt] MT FP & EDF Callbacks dispatched to OS worker threads, preemptive (no wait set)
@@
Liu et al. [liu2025rosrt] May, 2025 Humble https://github.com/sizheliu-unc/rclcpp
@@ 4.2 System-Level Enhancements
Optimizing Inter-Process Communication. Beyond CPU scheduling, researchers have proposed system-level enhancements targeting the performance and predictability of ROS 2’s inter-process communication (IPC), often focusing on limitations within the default DDS middleware or exploring alternatives. Kim et al. [kim2025cros] identified the lack of priority propagation across ROS 2’s multi-layered architecture, leading to priority inversion where low-priority DDS or kernel threads delay high-priority application messages. To address this, they developed the Cros-Rt framework, which explicitly aligns thread priorities across the application, middleware, and kernel layers to ensure consistent communication across layers that uphold priority ordering. Orthogonally, Ishikawa et al. [ishikawaaso2025ros2agnocastsupporting] proposed the ROS 2 Agnocast framework to optimize the data transfer efficiency within ROS 2 by using zero-copy shared memory for IPC. They observed that current applications are often bottlenecked by data serialization and copying overheads in DDS, especially for large messages, but also identified key limitations in existing zero-copy DDS implementations, such as the inability to handle unsized message types (e.g., images or point clouds). They introduce a novel publish/subscribe solution through Agnocast using shared memory and pointer-swapping that supports unsized message types, enabling true zero-copy IPC and further enhancing communication efficiency through latency reduction and throughput improvement. Specific middleware protocols for large data objects have been investigated also by Peeck et al. [Peeck2021, Peeck2023] in the context of the DDS. Luo et al. [luo2025flexible] proposed a multi-stage approach for enabling efficient zero-copy inter-process communication in ROS 2, with a focus on supporting dynamically structured data fields. The approach builds upon the mini memory management system component, allowing multiple processes to access the same message instance from a shared memory space, and the message propagation is adapted, which ensures compatibility with the existing ROS 2 communication framework.
Enhancing GPU and Accelerator Management. ROS 2 serves as a powerful resource manager for CPU tasks, but it lacks built-in support for GPU resource management. Furthermore, the vast majority of the discussed works focus on CPU scheduling only and neglect the presence of hardware accelerators. Nevertheless, many critical functionalities in autonomous systems, such as perception and intelligent control, require intensive GPU computation. However, under default ROS 2, these components must submit their GPU workloads independently, without coordination.